# When Gaussian Meets Surfel: Ultra-fast High-fidelity Radiance Field Rendering

### v1

@@ Abstract.
We introduce Gaussian-enhanced Surfels (GESs), a bi-scale representation for radiance field rendering, wherein a set of 2D opaque surfels with view-dependent colors represent the coarse-scale geometry and appearance of scenes, and a few 3D Gaussians surrounding the surfels supplement fine-scale appearance details. The rendering with GESs consists of two passes – surfels are first rasterized through a standard graphics pipeline to produce depth and color maps, and then Gaussians are splatted with depth testing and color accumulation on each pixel order independently. The optimization of GESs from multi-view images is performed through an elaborate coarse-to-fine procedure, faithfully capturing rich scene appearance. The entirely sorting-free rendering of GESs not only achieves very fast rates, but also produces view-consistent images, successfully avoiding popping artifacts under view changes. The basic GES representation can be easily extended to achieve anti-aliasing in rendering (Mip-GES), boosted rendering speeds (Speedy-GES) and compact storage (Compact-GES), and reconstruct better scene geometries by replacing 3D Gaussians with 2D Gaussians (2D-GES). Experimental results show that GESs advance the state-of-the-arts as a compelling representation for ultra-fast high-fidelity radiance field rendering.
@@ 1. Introduction
Free-view synthesis of 3D scenes from multi-view images has been a long-standing research topic for decades. Starting from the seminal work of light fields (Levoy and Hanrahan, 1996; Gortler et al., 1996), various scene representations have been developed for this task, with remarkable advances in recent years, e.g., Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023).
@@
In this paper, we introduce Gaussian-enhanced Surfels (GESs), a bi-scale representation for radiance field rendering. At the coarse scale, a set of 2D opaque surfels with view-dependent colors constitute an approximation of the surface radiance field of the scene, called the surfel radiance field. Each surfel is a 2D opaque ellipse associated with geometry attributes of position, rotation and scaling, and appearance attributes of spherical harmonics (SH) coefficients. At the fine scale, a few 3D Gaussians surrounding the coarse-scale surfels form a volumetric radiance field to supplement the scene appearance not represented well by the surfel radiance field. Each Gaussian has the same geometry and appearance attributes as in 3DGS (Kerbl et al., 2023). Our key observation is that the surfel radiance field can capture a major portion of the scene appearance (see Fig. 2, left), while the Gaussian radiance field effectively enhances the surfel radiance field with fine details (see Fig. 2, middle). Combining radiance renderings from both coarse and fine scales, the GES radiance field is able to synthesize high-quality images (see Fig. 2, right), competitive with state-of-the-art methods.
@@
The basic GES representation can be easily extended to further enhance its capability by incorporating recent improvements of the vanilla 3DGS method. By combining the filtering algorithm for Gaussian splatting (Yu et al., 2024a) with the standard multi sampling anti-aliasing (MSAA) for surfel rasterization, we are able to significantly reduce aliasing artifacts in renderred images (Mip-GES). The Hessian pruning score (Hanson et al., 2024) can be employed to largely reduce the number of Gaussians, further boosting our rendering speed (Speedy-GES). Following (Lee et al., 2024), we can replace the SH coefficients of both surfels and Gaussians by querying colors from a hash grid, and quantize the scaling and rotation of surfels and Gaussians, obtaining a compact storage of GES (Compact-GES). We can also replace the 3D Gaussians with 2D Gaussians (Huang et al., 2024) to reconstruct better scene geometries (2D-GES).
@@ 2.1. Traditional Scene Reconstruction and Rendering
Traditional approaches for novel view synthesis construct light fields from densely sampled images (Gortler et al., 1996; Levoy and Hanrahan, 1996; Buehler et al., 2001; Davis et al., 2012). With the development of Structure-from-Motion (SfM) (Snavely et al., 2006) and multi-view stereo (MVS) (Goesele et al., 2007), full 3D reconstruction is enabled from a collection of photos, based on which different view synthesis methods are proposed (Chaurasia et al., 2013; Hedman et al., 2018a; Kopanas et al., 2021). Among these prior works, the surface light field (Wood et al., 2000; Chen et al., 2018) is the most relevant representation to ours, which represents the radiance of rays originating from any points on the surface in any directions. The construction of surface light fields requires high-resolution geometry and dense sets of images. Differently, our GES representation uses surfels with view-dependent colors to approximate the surface light field and employs Gaussians to supplement the scene appearance not captured well by surfels. We do not require high-resolution geometry either.
@@ 2.2. Neural Radiance Fields
Different neural rendering approaches have been studied for novel view synthesis (Thies et al., 2019; Lombardi et al., 2019, 2021; Zhou et al., 2018; Wizadwongsa et al., 2021; Riegler and Koltun, 2020). Among the many successful approaches, NeRFs (Mildenhall et al., 2021) in particular have achieved remarkable success and have produced an explosion of follow-up works (Barron et al., 2021, 2022a; Tancik et al., 2022; Turki et al., 2022; Sun et al., 2022; Fridovich-Keil et al., 2022; Garbin et al., 2021; Yu et al., 2021; Takikawa et al., 2022; Reiser et al., 2021). For example, to accelerate the training and rendering of NeRFs, InstantNGP (Müller et al., 2022) replaces the deep MLP by a shallow MLP with the multiresolution hash encoding as input, and can be trained in a few minutes. Adaptive Shells (Wang et al., 2023) extract two meshes from SDFs as scene boundaries. During rendering, it only samples points between ray-mesh intersections to significantly reduce the number of volumerric samples. Quadrature Fields (Sharma et al., 2024) extract meshes from a quadrature field built with the help of NeRFs and bake view-dependent colors obtained from NeRFs as compressed spherical Gaussians stored in textures, after which ray tracing is performed to get all ray-mesh intersections as samples for volume rendering.
@@ 2.3. Point-based Rendering
Point-based rendering has been explored for decades for rendering disconnected and unstructured point clouds. To achieve high rendering quality, a common solution is to render points as larger primitives such as circular or elliptic discs, ellipsoids, or surfels (Botsch et al., 2005; Pfister et al., 2000; Ren et al., 2002a; Zwicker et al., 2001, 2004). With the development of differentiable point-based rendering (Wiles et al., 2020; Yifan et al., 2019), point based rendering techniques are also adopted for novel view synthesis (Aliev et al., 2020; Rückert et al., 2022; Kopanas et al., 2021; Lassner and Zollhöfer, 2021; Kopanas et al., 2022). The most recent work (Zhang et al., 2022) in this category presents point-based radiance fields, where each point is associated with the properties of position and SH coefficients, and has no other geometry properties (e.g., size, rotation or normal). The points are rendered with point splatting, where an image-space Gaussian radial basis function is employed to compute the alpha value of each point on each pixel. The points are then sorted by z-distance and the image is rendered via alpha blending from front to back. The method is designed for foreground scene components and requires masks for initialization, and it is unclear how it can scale to general scenes. Point-NeRF (2022) also uses points to represent radiance fields, but it still requires MLPs and volumetric ray-marching, and hence cannot achieve real-time rendering.
@@
3DGS (Kerbl et al., 2023) represents radiance fields as sparsely distributed 3D Gaussians with view-dependent colors. Due to the exceptionally fast speed and high quality in view synthesis, it has inspired a great amount of follow-up research (Moenne-Loccoz et al., 2024a; Hahlbohm et al., 2024; Kerbl et al., 2024; Wu et al., 2024a; Yu et al., 2024a; Papantonakis et al., 2024; Fan et al., 2024; Lyu et al., 2024; Yu et al., 2024c). (Chen and Wang, 2024) and (Wu et al., 2024b) provide a comprehensive coverage of recent advancements in this field.
@@
Among these works, 2DGS (Huang et al., 2024) and Gaussian surfels (Dai et al., 2024) adopt 2D Gaussians instead of 3D Gaussians to represent radiance fields. By introducing a perspective-accurate 2D splatting process and incorporating depth distortion and normal consistency terms, 2DGS can optimize 2D Gaussians to be distributed more closely around the surface, hence producing more geometrically accurate radiance fields. SolidGS (Shen et al., 2024) modifies the 2D Gaussian kernel function to create a larger opaque region for improved view consistency. Geometry Field Splatting (Jiang et al., 2024) uses 2D Gaussians to represent a geometry field and converts it into a density field for volumetric rendering. Nevertheless, all of these methods are still volumetric representations of radiance fields – its rendering requires sorting and -blending of all Gaussians along the ray for each pixel as in 3DGS.
@@
It is known that 3DGS suffers from the popping artifacts during camera movements, because it approximates the accurate per-pixel depth sorting with a tile-based global sorting of Gaussian center depths. StopThePop (Radl et al., 2024) proposes a hierarchical sorting strategy to alleviate the popping artifacts Hahlbohm et al. (2024) use hybrid transparency (Wyman, 2016) to approximate the accurate blending per-pixel. Both methods cannot guarantee the per-pixel ordering is fully correct, and thus could produce popping artifacts in novel view synthesis (see Fig. 12 and the supplementary video). Hou et al. (2024a) present a sorting-free method for 3DGS by approximating alpha blending with weighted sums, thereby eliminating the popping artifacts. However, because for each pixel it calculates the weighted sum of all Gaussians along the ray, this method will produce color leakage artifacts of occluded objects.
@@
Many methods have been proposed to accelerate 3DGS rendering and reduce storage by pruning redundant Gaussians (Fan et al., 2024; Lee et al., 2024; Niemeyer et al., 2024; Hanson et al., 2024). While very fast rendering can be achieved, the excessive pruning of Gaussians tends to lose appearance details. More importantly, as the remaining Gaussians contribute more significantly to the color, the popping artifact is exacerbated. Other methods improve the rasterization pipeline (Wang et al., 2024; Feng et al., 2024), such as reducing the number of covered tiles of screen space Gaussians. However, as sorting still exists as a computational bottleneck, their rendering speed is slower than our method.
@@
RTG-SLAM (Peng et al., 2024) presents a real-time 3D reconstruction system with an RGBD camera for large-scale environments using 3D Gaussian splatting. It uses opaque Gaussians to fit the geometry and dominant colors and nearly-transparent Gaussians to fit residual colors. The opacity value of the opaque Gaussians is 0.99 in the center and decays outwards with the Gaussian function, while the opacity value of our opaque surfels is 1 within the whole surfel range. The rendering of RTG-SLAM is the same as 3DGS.
@@
Our surfel optimization is performed by utilizing the opacity modulating parameter to gradually evolve the translucent surfels into opaque ones. This idea is related to differentiable mesh rasterizer methods. They typically leverage the subpixel antialiasing operation (Laine et al., 2020) or convert the mesh into a collection of 3D Gaussians (Rhodin et al., 2015) or build smooth probability maps of each triangle (Liu et al., 2019) to get the gradients of image loss related to vertex positions. Our design is different, which gradually optimizes translucent surfels into opaque ones using the opacity modulating parameter.
@@ 3. Gaussian-enhanced Surfels
The 2D surfel is defined as a 2D unit circular disc on the XY-plane in its local coordinate, associated with a set of properties as , where is the surfel center position, is the rotation quaternion, is the anisotropic scaling, and is the spherical harmonics coefficients representing view-dependent colors. This 2D surfel forumla is similar to the 2D Gaussian of 2DGS (Huang et al., 2024), except that our surfels are fully opaque. The surfel is transformed to the world space by sequentially applying the scaling, rotation and translation with . The color of the whole surfel is the same as the SH color in the direction from the surfel center to the camera position as
@@
where is the indicator function, is the Gaussian center depth, is the pixel depth obtained from the depth map , is the projected covariance matrix determined by the rotation and scaling, is a small positive value introduced to prevent the 3D Gaussians distributed close to the surface from being wrongly truncated. is the projected Gaussian center, and is the Gaussian color in the current view. is clamped to zero if , as in 3DGS (Kerbl et al., 2023).
@@ 4. Coarse-to-Fine Optimization
Given a set of multi-view images of a static scene, whose corresponding cameras are calibrated by SfM (Schönberger and Frahm, 2016), our goal is to optimize a set of 2D surfels and 3D Gaussians faithfully representing the scene for free-view synthesis. The optimization is performed using a coarse-to-fine method. The coarse stage optimizes the opaque surfels to reconstruct the coarse-scale geometry and appearance. In the fine stage, the surfel geometry properties (i.e., and ) are fixed, and the 3D Gaussians are added and are jointly optimized together with the surfel SH coefficients to reconstruct rich appearance details.
@@ 4.1. Surfel Optimization
where is clamped to 0 when . The rasterization of translucent surfels is the same as in 2DGS (Huang et al., 2024). We use the opacity modulating parameter to establish the connection between the translucent surfel and the opaque surfel. When , the surfel is a 2D Gaussian. When increases, the Gaussian gradually becomes more opaque from the center outward. When equals 255, the opacity is 1 within the range where , and 0 outside this range. This results in a circular disc with a radius of where the opacity is 1 uniformly, as illustrated in Fig. 4.
During optimization, the outer ring of the surfel is semi-transparent with Gaussian distributed opacity, which allows color-based gradients to be backpropagated to the surfel geometry parameters. The optimization starts with , which means each surfel is a 2D Gaussians. In the early iterations, the large semi-transparent regions provide effective gradients for optimizing the position and shape of the surfels. We increase gradually in the later iterations, and the semi-transparent regions shrink accordingly, with the surfel geometry eventually stabilized until the translucent surfel becomes a fully opaque one. To obtain more gradients in semi-transparent regions and achieve anti-aliasing, we use supersampling once the values of all surfels are increased over 30.
@@
The loss function includes an image loss and a D-SSIM loss as in (Kerbl et al., 2023):
where in our implementation. In the first 10K iterations, we use the same strategy as in (Yu et al., 2024a) for surfel densification and pruning. At the 10K-th iteration, we discard the surfels with (but record the position of these surfels for subsequent 3D Gaussian initialization), increase of the remaining surfels to no smaller than 30, keep from optimization, and disable the densification and pruning. In the 18K-th and 19K-th iteration, we increase of each surfel to no smaller than 60 and 90, respectively. In the 20K-th iteration, we set for all surfels and terminate surfel geometry optimization.

### v2

@@ Abstract.
We introduce Gaussian-enhanced Surfels (GESs), a bi-scale representation for radiance field rendering, wherein a set of 2D opaque surfels with view-dependent colors represent the coarse-scale geometry and appearance of scenes, and a few 3D Gaussians surrounding the surfels supplement fine-scale appearance details. The rendering with GESs consists of two passes – surfels are first rasterized through a standard graphics pipeline to produce depth and color maps, and then Gaussians are splatted with depth testing and color accumulation on each pixel order independently. The optimization of GESs from multi-view images is performed through an elaborate coarse-to-fine procedure, faithfully capturing rich scene appearance. The entirely sorting-free rendering of GESs not only achieves very fast rates, but also produces view-consistent images, successfully avoiding popping artifacts under view changes. The basic GES representation can be easily extended to achieve anti-aliasing in rendering (Mip-GES), boosted rendering speeds (Speedy-GES) and compact storage (Compact-GES), and reconstruct better scene geometries by replacing 3D Gaussians with 2D Gaussians (2D-GES). Experimental results show that GESs advance the state-of-the-arts as a compelling representation for ultra-fast high-fidelity radiance field rendering. Code and the interactive web viewer are available online.
@@ 1. Introduction
Free-view synthesis of 3D scenes from multi-view images has been a long-standing research topic for decades. Starting from the seminal work of light fields (lightfield1996; lumigraph1996), various scene representations have been developed for this task, with remarkable advances in recent years, e.g., Neural Radiance Fields (NeRFs) (nerf) and 3D Gaussian Splatting (3DGS) (3DGS).
@@
In this paper, we introduce Gaussian-enhanced Surfels (GESs), a bi-scale representation for radiance field rendering. At the coarse scale, a set of 2D opaque surfels with view-dependent colors constitute an approximation of the surface radiance field of the scene, called the surfel radiance field. Each surfel is a 2D opaque ellipse associated with geometry attributes of position, rotation and scaling, and appearance attributes of spherical harmonics (SH) coefficients. At the fine scale, a few 3D Gaussians surrounding the coarse-scale surfels form a volumetric radiance field to supplement the scene appearance not represented well by the surfel radiance field. Each Gaussian has the same geometry and appearance attributes as in 3DGS (3DGS). Our key observation is that the surfel radiance field can capture a major portion of the scene appearance (see Fig. 2, left), while the Gaussian radiance field effectively enhances the surfel radiance field with fine details (see Fig. 2, middle). Combining radiance renderings from both coarse and fine scales, the GES radiance field is able to synthesize high-quality images (see Fig. 2, right), competitive with state-of-the-art methods.
@@
The basic GES representation can be easily extended to further enhance its capability by incorporating recent improvements of the vanilla 3DGS method. By combining the filtering algorithm for Gaussian splatting (mipgs2024) with the standard multi sampling anti-aliasing (MSAA) for surfel rasterization, we are able to significantly reduce aliasing artifacts in renderred images (Mip-GES). The Hessian pruning score (hanson2024speedy) can be employed to largely reduce the number of Gaussians, further boosting our rendering speed (Speedy-GES). Following (lee2024compact), we can replace the SH coefficients of both surfels and Gaussians by querying colors from a hash grid, and quantize the scaling and rotation of surfels and Gaussians, obtaining a compact storage of GES (Compact-GES). We can also replace the 3D Gaussians with 2D Gaussians (2dgs2024) to reconstruct better scene geometries (2D-GES).
@@ 2.1. Traditional Scene Reconstruction and Rendering
Traditional approaches for novel view synthesis construct light fields from densely sampled images (lumigraph1996; lightfield1996; unstructlumi2001; davis2012unstructured). With the development of Structure-from-Motion (SfM) (sfm2006) and multi-view stereo (MVS) (mvs2007), full 3D reconstruction is enabled from a collection of photos, based on which different view synthesis methods are proposed (dsin2013; deepblending2018; pnrvo2021). Among these prior works, the surface light field (surfacelf2000; AnpeiChen2018) is the most relevant representation to ours, which represents the radiance of rays originating from any points on the surface in any directions. The construction of surface light fields requires high-resolution geometry and dense sets of images. Differently, our GES representation uses surfels with view-dependent colors to approximate the surface light field and employs Gaussians to supplement the scene appearance not captured well by surfels. We do not require high-resolution geometry either.
@@ 2.2. Neural Radiance Fields
Different neural rendering approaches have been studied for novel view synthesis (neuraltexture2019; neuralvolume2019; MVP2021; MPI2018; NeX2021; fvs2020). Among the many successful approaches, NeRFs (nerf2021) in particular have achieved remarkable success and have produced an explosion of follow-up works (mipnerf2021; mipnerf360; blocknerf2022; MegaNerf2022; DVGO2022; Plenoxels2022; fastnerf2021; PlenOctrees2021; VBNF2022; KiloNeRF2021). For example, to accelerate the training and rendering of NeRFs, InstantNGP (InstantNGP2022) replaces the deep MLP by a shallow MLP with the multiresolution hash encoding as input, and can be trained in a few minutes. Adaptive Shells (adaptiveshells2023) extract two meshes from SDFs as scene boundaries. During rendering, it only samples points between ray-mesh intersections to significantly reduce the number of volumerric samples. Quadrature Fields (sharma2024volumetric) extract meshes from a quadrature field built with the help of NeRFs and bake view-dependent colors obtained from NeRFs as compressed spherical Gaussians stored in textures, after which ray tracing is performed to get all ray-mesh intersections as samples for volume rendering.
@@ 2.3. Point-based Rendering
Point-based rendering has been explored for decades for rendering disconnected and unstructured point clouds. To achieve high rendering quality, a common solution is to render points as larger primitives such as circular or elliptic discs, ellipsoids, or surfels (hqss2005; surfel2000; OS_EWA_SS2002; SS2001; PSS2004). With the development of differentiable point-based rendering (SynSin2020; DSS2019), point based rendering techniques are also adopted for novel view synthesis (NPG2020; ADOP2022; pnrvo2021; Pulsar2021; catacaustics). The most recent work (DPRF2022) in this category presents point-based radiance fields, where each point is associated with the properties of position and SH coefficients, and has no other geometry properties (e.g., size, rotation or normal). The points are rendered with point splatting, where an image-space Gaussian radial basis function is employed to compute the alpha value of each point on each pixel. The points are then sorted by z-distance and the image is rendered via alpha blending from front to back. The method is designed for foreground scene components and requires masks for initialization, and it is unclear how it can scale to general scenes. Point-NeRF (pointnerf2022) also uses points to represent radiance fields, but it still requires MLPs and volumetric ray-marching, and hence cannot achieve real-time rendering.
@@
3DGS (3DGS) represents radiance fields as sparsely distributed 3D Gaussians with view-dependent colors. Due to the exceptionally fast speed and high quality in view synthesis, it has inspired a great amount of follow-up research (3dgsraytracing2024; 3dgsht2024; hier3dgs2024; 4dgs2024; mipgs2024; reducemem3dgs2024; lightGaussian2023; 3DGSR2024; GOF2024). (3dgssurvey) and (3DGSsurvey-CVM2024) provide a comprehensive coverage of recent advancements in this field.
@@
Among these works, 2DGS (2dgs2024) and Gaussian surfels (dai2024high) adopt 2D Gaussians instead of 3D Gaussians to represent radiance fields. By introducing a perspective-accurate 2D splatting process and incorporating depth distortion and normal consistency terms, 2DGS can optimize 2D Gaussians to be distributed more closely around the surface, hence producing more geometrically accurate radiance fields. SolidGS (shen2024solidgs) modifies the 2D Gaussian kernel function to create a larger opaque region for improved view consistency. Geometry Field Splatting (jiang2024geometry) uses 2D Gaussians to represent a geometry field and converts it into a density field for volumetric rendering. Nevertheless, all of these methods are still volumetric representations of radiance fields – its rendering requires sorting and -blending of all Gaussians along the ray for each pixel as in 3DGS.
@@
It is known that 3DGS suffers from the popping artifacts during camera movements, because it approximates the accurate per-pixel depth sorting with a tile-based global sorting of Gaussian center depths. StopThePop (stopthepop2024) proposes a hierarchical sorting strategy to alleviate the popping artifacts Hahlbohm et al. (3dgsht2024) use hybrid transparency (hybridtrans2016) to approximate the accurate blending per-pixel. Both methods cannot guarantee the per-pixel ordering is fully correct, and thus could produce popping artifacts in novel view synthesis (see Fig. 12 and the supplementary video). Hou et al. (sortfree2024) present a sorting-free method for 3DGS by approximating alpha blending with weighted sums, thereby eliminating the popping artifacts. However, because for each pixel it calculates the weighted sum of all Gaussians along the ray, this method will produce color leakage artifacts of occluded objects.
@@
Many methods have been proposed to accelerate 3DGS rendering and reduce storage by pruning redundant Gaussians (lightGaussian2023; lee2024compact; niemeyer2024radsplat; hanson2024speedy). While very fast rendering can be achieved, the excessive pruning of Gaussians tends to lose appearance details. More importantly, as the remaining Gaussians contribute more significantly to the color, the popping artifact is exacerbated. Other methods improve the rasterization pipeline (adr2024; feng2024flashgs), such as reducing the number of covered tiles of screen space Gaussians. However, as sorting still exists as a computational bottleneck, their rendering speed is slower than our method.
@@
RTG-SLAM (RTGSLAM) presents a real-time 3D reconstruction system with an RGBD camera for large-scale environments using 3D Gaussian splatting. It uses opaque Gaussians to fit the geometry and dominant colors and nearly-transparent Gaussians to fit residual colors. The opacity value of the opaque Gaussians is 0.99 in the center and decays outwards with the Gaussian function, while the opacity value of our opaque surfels is 1 within the whole surfel range. The rendering of RTG-SLAM is the same as 3DGS.
@@
Our surfel optimization is performed by utilizing the opacity modulating parameter to gradually evolve the translucent surfels into opaque ones. This idea is related to differentiable mesh rasterizer methods. They typically leverage the subpixel antialiasing operation (Laine2020diffrast) or convert the mesh into a collection of 3D Gaussians (rhodin2015versatile) or build smooth probability maps of each triangle (softrasterizer19) to get the gradients of image loss related to vertex positions. Our design is different, which gradually optimizes translucent surfels into opaque ones using the opacity modulating parameter.
@@ 3. Gaussian-enhanced Surfels
The 2D surfel is defined as a 2D unit circular disc on the XY-plane in its local coordinate, associated with a set of properties as , where is the surfel center position, is the rotation quaternion, is the anisotropic scaling, and is the spherical harmonics coefficients representing view-dependent colors. This 2D surfel forumla is similar to the 2D Gaussian of 2DGS (2dgs2024), except that our surfels are fully opaque. The surfel is transformed to the world space by sequentially applying the scaling, rotation and translation with . The color of the whole surfel is the same as the SH color in the direction from the surfel center to the camera position as
@@
where is the indicator function, is the Gaussian center depth, is the pixel depth obtained from the depth map , is the projected covariance matrix determined by the rotation and scaling, is a small positive value introduced to prevent the 3D Gaussians distributed close to the surface from being wrongly truncated. is the projected Gaussian center, and is the Gaussian color in the current view. is clamped to zero if , as in 3DGS (3DGS).
@@ 4. Coarse-to-Fine Optimization
Given a set of multi-view images of a static scene, whose corresponding cameras are calibrated by SfM (sfm2016), our goal is to optimize a set of 2D surfels and 3D Gaussians faithfully representing the scene for free-view synthesis. The optimization is performed using a coarse-to-fine method. The coarse stage optimizes the opaque surfels to reconstruct the coarse-scale geometry and appearance. In the fine stage, the surfel geometry properties (i.e., and ) are fixed, and the 3D Gaussians are added and are jointly optimized together with the surfel SH coefficients to reconstruct rich appearance details.
@@ 4.1. Surfel Optimization
where is clamped to 0 when . The rasterization of translucent surfels is the same as in 2DGS (2dgs2024). We use the opacity modulating parameter to establish the connection between the translucent surfel and the opaque surfel. When , the surfel is a 2D Gaussian. When increases, the Gaussian gradually becomes more opaque from the center outward. When equals 255, the opacity is 1 within the range where , and 0 outside this range. This results in a circular disc with a radius of where the opacity is 1 uniformly, as illustrated in Fig. 4.
During optimization, the outer ring of the surfel is semi-transparent with Gaussian distributed opacity, which allows color-based gradients to be backpropagated to the surfel geometry parameters. The optimization starts with , which means each surfel is a 2D Gaussian. In the early iterations, the large semi-transparent regions provide effective gradients for optimizing the position and shape of the surfels. We increase gradually in the later iterations, and the semi-transparent regions shrink accordingly, with the surfel geometry eventually stabilized until the translucent surfel becomes a fully opaque one. To obtain more gradients in semi-transparent regions and achieve anti-aliasing, we use supersampling once the values of all surfels are increased over 30.
@@
The loss function includes an image loss and a D-SSIM loss as in (3DGS):
where in our implementation. In the first 10K iterations, we use the same strategy as in (mipgs2024) for surfel densification and pruning. At the 10K-th iteration, we discard the surfels with (but record the position of these surfels for subsequent 3D Gaussian initialization), increase of the remaining surfels to no smaller than 30, keep from optimization, and disable the densification and pruning. In the 18K-th and 19K-th iteration, we increase of each surfel to no smaller than 60 and 90, respectively. In the 20K-th iteration, we set for all surfels and terminate surfel geometry optimization.